Assuming you are using 1.0 Å wavelength X-rays: a. What is the scattering angle for reflections at 3.0 Å resolution? b. 2.0 Å resolution? c. 1.5 Å resolution? Recall that the scattering angle is 2x the angle θ in Bragg’s Law.
Answer #1

We know that

n\lambda = 2d\sin \theta

For n = 1

\lambda = 2d\sin \theta

\sin \theta =\frac{\lambda}{ 2d}

\theta =\sin^{-1}\left (\frac{\lambda}{ 2d} \right )

2x = \theta =\sin^{-1}\left (\frac{\lambda}{ 2d} \right )

x = \frac{\theta}{2} =\frac{1}{2}\sin^{-1}\left (\frac{\lambda}{ 2d} \right )

Part A) \lambda = 1.0 \AA , d = 3.0 \AA

x = \frac{1}{2}\sin^{-1}\left (\frac{1.0\AA }{ 2\times 3.0\AA } \right ) = 4.8^{0}

Part B) \lambda = 1.0 \AA, d = 2.0 \AA

x = \frac{1}{2}\sin^{-1}\left (\frac{1.0\AA }{ 2\times 2.0\AA } \right ) = 7.24^{0}

Part C) \lambda = 1.0 \AA , d = 1.5 \AA

x = \frac{1}{2}\sin^{-1}\left (\frac{1.0\AA }{ 2\times 1.5\AA } \right ) = 9.74^{0}